OpenStudy (austine736):

What is the answer to this problem, t(8 – 3) = 2 • 8 – 2 • 3

TheSmartOne (thesmartone):

So, did anywhere did they say that t = 2?

OpenStudy (mathwizzard3):

what do you think you should do first?

TheSmartOne (thesmartone):

So, did they say anywhere that t = 2? **

OpenStudy (austine736):

yes

TheSmartOne (thesmartone):

Well, first since t = 2, let's plug it in: t(8-3) = 2(8-3) Can you tell me what 8 - 3= ?

OpenStudy (austine736):

5

TheSmartOne (thesmartone):

ok, now we have: 2(8 - 3) = 2(3) What is 2 x 3 = ?

TheSmartOne (thesmartone):

@austine736

OpenStudy (austine736):

6

TheSmartOne (thesmartone):

and so that is your final answer :)

OpenStudy (austine736):

thanks

OpenStudy (mathmale):

Austine: Please always add the instructions that came with a given mathematical expression. In this case, the instructions were "Solve the following expression for t."